# Social Media Should Feel Like Minecraft, Not Instagram: 3D Gamer Youth Visions for Meaningful Social Connections through Fictional Inquiry

### v2

@@ 4.5.2. Making Social Media About (Meaningful) Connection, Not (Passive) Consumption
Beyond how users navigate content, participants also emphasized the importance of shifting the platform’s core purpose from endless consumption toward meaningful social connection. Participants expressed frustration with how current platforms promote mindless, passive consumption, and described their hopes for SMH to instead foster more meaningful, intentional connections through active social engagement. They described frustration with how easily time slips away on platforms like Instagram and TikTok without leaving them feeling unfulfilled. P10 noted that social media often feels “meaningless” when they “didn’t learn something,” and P18 reflected on the regret of “mindlessly scrolling.” P14 similarly described spending time on Instagram Reels while “literally just laying in [their] bed” as “wasting [their] time.”
A major contributor to this feeling was the infinite scrolling design of MSM. P11 explained that “there’s always more content on social media,” which fuels the fear of missing out and the urge to keep scrolling. P14 critiqued TikTok for exposing users to “clickbaity” and often low-quality content that “doesn’t stimulate [users] at all” and “messes with [their] attention span.”
In contrast, participants described active, social engagement, even if casual, as far more rewarding. P16 shared that “anytime not scrolling” but instead “talking to people” feels worthwhile, “even if it’s just light conversation.” P14 echoed this, noting that meaningful use comes when they are “actually talking to someone,” especially to stay connected with people “not in close proximity.” They reflected that while users might spend more time on a platform like SMH, that time would feel worthwhile because “at the end of the day, you’re connecting with a person that you can’t otherwise connect with… and that makes it more valuable in and of itself” compared to passively watching “50-second TikTok videos of somebody making slime for like an hour.”
Participants also contrasted passive consumption with active participation in interactive environments. P10 explained that while doomscrolling feels unproductive, activities like playing Minecraft feel different because they offer “progress” that can be “saved and come back to.” P06 echoed this sentiment, noting that they “don’t feel bad… most of the time [they] use Roblox,” because it involves “playing games with friends” and “customizing spaces.”
Reflecting on SMH’s potential to create this kind of active, engaging experience, P09 acknowledged that “time spent might not be mitigated,” but emphasized that “the way you’re spending it” would make the difference. They explained that SMH would “feel a lot more interactive instead of just taking in content, and more and more and more content,” suggesting that meaningful engagement, rather than passive consumption, was what ultimately mattered for them.
@@ 4.6. Playfulness and Whimsy
Participants described SMH as a platform that feels youthful, playful, and deeply immersive, offering not just entertainment but opportunities for creative self-expression, social bonding, and emotional well-being. By embracing whimsy and delight, SMH was seen as making social media feel more personal, less pressured, and more meaningful.
@@ 4.6.1. Embracing Creative Joy and Playful Self-Expression
Participants imagined SMH as a platform that transforms social media into a space for creative joy and playful self-expression, offering users the freedom to experiment with identity, explore personal interests, and connect with others through creativity. Several participants described how SMH offered far more than static profiles or simple feeds. P06 reflected that “everyone likes sharing parts of themselves,” appreciating how the platform allows users to shape not just what they say, but how they appear and interact with others. P11 imagined a “dream social media world” where users could fully redesign their profiles, “maybe their avatar in the middle, maybe their posts on either side, maybe a bio below”, allowing them to “express their creativity and, like, yourself” in new ways not possible on conventional platforms.
For some, this act of personal customization felt like a meaningful process of identity exploration. P01 described the deeper satisfaction in “just that process of creating it, and finding out who you are through that process.” P21 echoed this sentiment, describing SMH as feeling “more like a game” because of the “creative freedom” it provides, making the experience feel “more fictional than real” in a way that encourages experimentation and self-expression. P22 added that SMH offered something they hadn’t found anywhere else, explaining, “I can really be creative here… I can make something that I haven’t made anywhere else, you know. That’s where the fun aspect comes in, really, for me.”
This spirit of creativity extends beyond the individual to the wider environment and community. Participants described the excitement of exploring what others had created, likening it to visiting realms in Minecraft where you “get to see how like everyone puts in their own effort into it… Whoa! They built that, that’s so cool” (P06). SMH felt alive and engaging whether users were socializing or spending time alone: “it’s not just like, Oh, the group chats are only fun when there’s other people here. No, it’s fun when we’re by ourselves, too, we get to do our own things” (P06). The interactive environment itself, where “you’re not just sitting there chatting, you’re also walking around the space… you’re interacting with whatever else is in that area” (P06), offered more to do than conventional platforms. Whether through building, exploring, or engaging with shared content, SMH transformed social interaction into something playful, creative, and deeply engaging.
@@ 4.6.2. Designing Engaging and Playful Privacy Mechanisms
Participants imagined privacy not as a tedious and confusing task but as part of the playful experience, envisioning interactive, game-like mechanisms that turn boundary-setting into something fun, discoverable, and socially meaningful. P11 envisioned systems where users could control access to spaces through gestures like a “secret knock” or hidden objects that revealed “secret passages” accessible only to trusted friends. Others suggested knowledge-based gates, such as “secret handshakes” (P21) shared among friends. Drawing inspiration from the Marauder’s Map, P02 imagined privacy controls activated by “spells”, unlocking spaces only when correctly summoned.
These features were described not just as functional but as adding delight and novelty to social interaction. P05 compared them to “little Easter eggs,” remarking that “there’s that little kid part of everyone” that makes such features enjoyable. They contrasted these with conventional privacy settings on mainstream social media, which they described as “a wall of settings” that often feel confusing or uninviting.
@@ 4.6.3. Augmenting Everyday Interactions with Fantastical Interfaces and Environments
Participants imagined SMH using magical and fantastical technologies to transform everyday digital interactions, making them feel more immersive, accessible, and emotionally uplifting, even for those with limited real-world opportunities. P01 questioned the need for phones at all, asking, “If you’re already carrying around like a wand everywhere with you, why would you also want to carry a phone?” They envisioned wands projecting “holographic screens, like Jarvis in Iron Man,” or activating “panels” and “big screens” through spells. Others suggested alternative magical devices. P15 preferred VR glasses, describing them as “more natural,” “easier,” and “more fun.” Participants also imagined objects like a “makeup mirror” that could serve as a portal, a “watch” that immerses users into SMH, or a “mirror glass” for video calls or summoning friends with spells like accio.
@@ 4.7. Expanding Horizons
Given the speculative nature of the FI co-design workshop, SMH embodies the youth’s aspirations and values for what interactions on social media platforms could embody. Therefore, while they acknowledged potential challenges, such as the platform being over-immersive or distracting, they overwhelmingly viewed these downsides as worthwhile trade-offs for its many benefits. Overall, participants described SMH as an inspiring alternative or expansion of existing social media, offering opportunities for self-expression and the ability to connect in ways that felt deeply personal and intentional. They saw SMH not just as a better platform, but as an inspiring redefinition of social media’s purpose and potential.
@@ 4.7.1. Breaking Free from Platform Monotony with Novel Interactions
Participants saw SMH as a platform that offers better user experiences, departs from the repetitive structures of existing platforms, and inspires hope for what social media can become. For example, P06 shared how SMH combines several desirable aspects of social media in a cohesive way, creating an environment focused on meaningful interactions: “You get friends who you can talk to, keep up with, and stay connected to. And you get things you can customize. You kind of get the whole package there. I kind of like all of those here.”
Participants also appreciated how SMH offered a break from the monotony of existing platforms. P11 pointed out that most social media platforms share a common “standard” with repetitive features like “feed, messages, settings,… profile.” In contrast, SMH introduced new types of interactions that deviated from familiar patterns. P02 described SMH as “a combination of Discord and BeReal” that brings together casual sharing and topic-based discussions. Similarly, P17 envisioned their SMH as “more interactive, kind of like if Discord and Instagram were combined,” where they could enjoy the interactions of Discord along with profiles and the ability to post content like on Instagram. P22 highlighted that on other platforms, there is limited opportunity for customization or decoration, explaining, “I just have a profile picture.” However, on SMH, they could “really be creative” and “make something that [they] haven’t made anywhere else,” which made the experience feel “novel,” like a “breath of fresh air.”
For some participants, SMH’s novelty lay in giving themselves permission to prioritize enjoyment over seriousness. P14 described initially defaulting to the idea of a more “mature” 2D version, but later embracing the playful, immersive 3D version as more personally meaningful and fun. For example, P14 initially leaned toward preferring a 2D version of the platform but shifted their preference to the 3D version, reasoning:
@@
P22 described SMH as “the coolest,” “childlike,” “less generic,” and “a lot more fun” compared to other platforms. P06 echoed this, observing that “the physical space” on SMH felt “a lot more fantastical… whimsical, fun” in contrast to “traditional ones.” For some, the platform also evokes nostalgia and happiness. P15 remarked, “Having this app would actually make them feel happy… with nostalgia.” They anticipated that SMH would be “a super big hit” with teenagers due to its playful and engaging nature.
@@ 4.7.2. Inspiring Hope and Expanding the Scope of Social Media Interactions
Through the FI process, participants realized they were not limited by the structures of existing platforms. Instead, they felt empowered to imagine social technologies that center joy, creativity, and authentic human connection, qualities they had once believed were beyond the reach of digital spaces. P10 pointed out that overcoming the design fixation around social media is crucial:
@@ 5. Discussion
Figure 2 captures how participants described their vision for ideal social media as something that pulls benefits from both types of platforms they already use. On one side are social networking and broadcasting platforms like Instagram, Facebook, and TikTok, which make it easy to stay socially visible but often feel shallow or performative. On the other side are platforms like FaceTime, Minecraft, Discord, and Roblox, where youth experience more organic and engaging connections, but where social interaction is often secondary to other goals like gaming or productivity. Our participants imagined “Social Media at Hogwarts (SMH)” as a platform that blends these strengths, creating a space where meaningful social connection effectively surfaces not just a peripheral effect, but a key focus of the platform.
@@ 5.2. Spatial Integrity for Meaningful Social Connections in Online Environments
5.3. Overcoming Cognitive Fixation and Doom Narratives Through Fictional Inquiry
However, taking social media away from young people or framing it as inherently harmful oversimplifies a much more complex reality. Despite public concern about social media’s role in youth mental health, the causal evidence remains inconclusive, and youth themselves report a mix of positive and negative experiences (Vogels, 2023). At the same time, research shows that fear-based policies and alarmist narratives can do harm of their own, amplifying anxiety, mistrust, and feelings of helplessness among youth, rather than supporting constructive engagement (Kim et al., 2024a; Lee and Hancock, 2024). This misalignment suggests that the solution lies not in rejecting social media altogether, but in asking what these platforms could become if they were designed to better reflect the kinds of social connection that youth actually seek. Rather than accepting the limitations of existing platforms as inevitable, there is an opportunity to rethink their underlying priorities and architectures.
Our interviews reveal how mainstream social media’s design fixations (Jansson and Smith, 1991) have shaped youth’s perceptions of what social platforms can be. Youth have come to equate social media with certain limiting characteristics: endless content streams, superficial interactions, and 2D interfaces focused on broadcasting posts. Interestingly, while they report meaningful social connections and positive experiences in other online social spaces like Minecraft or Discord, they view mainstream social media as inherently tied to emotionally disconnected experiences and superficial interactions. This cognitive fixation not only overlooks how social features can be designed differently—as demonstrated by their own designs—but also reinforces negative perceptions of social media that prior literature suggests can impact psychological well-being and quality of life (Lee and Hancock, 2024; Kim et al., 2024a).
To help youth envision beyond these constraints, we leveraged the FI method, immersing participants in a magical world setting. The Hogwarts-inspired context provided fresh metaphors for exploring social spaces, privacy, and connection. Unlike traditional design workshops where youth typically propose incremental improvements to existing platforms, these sessions enabled participants to imagine entirely new possibilities for meaningful connection. More importantly, the workshops helped participants expand their understanding of what social platforms could be, empowering them to articulate desires they might have previously dismissed as impossible within conventional social media paradigms. This study demonstrates how design methods can offer a generative path toward hope in domains where narratives of doom are prevalent.
@@ 5.4. Limitation and Future Work
Our study has several key limitations to consider. First, all participants had prior experience with 3D gaming platforms, so the findings do not inherently generalize to youth who are not familiar with such environments. Second, although we were cognizant of how fixed perceptions around “social media” might constrain ideation, our continued use of this terminology throughout the interviews may have inadvertently reinforced existing mental models rather than encouraging completely novel conceptualizations. Third, while the Harry Potter narrative framework successfully transported participants into a magical design space, our specific scenario about a Muggle-born student introducing social technology sometimes led participants to focus on wizard-Muggle divisions rather than the intended focus on connection across distance. This suggests opportunities to refine the narrative approach.
@@ 6. Conclusion
This study explored how 3D gamer youth reimagine social media when given the freedom to think beyond the limits of today’s platforms. Through the fictional world of Hogwarts, participants envisioned social experiences that felt spatial, relational, and emotionally meaningful, contrasting the passive, feed-based consumption that defines mainstream social media.
Their visions called for platforms that feel more like 3D games than archetypal ones like Instagram, emphasizing intentional navigation, ambient social presence, customizable spaces, and playful privacy mechanisms that adapt to the complexity of real relationships and incorporate elements of playfulness and whimsy.
Building on these insights, we introduced the concept of spatial integrity, a set of foundational design principles that make such visions possible. While spatial integrity alone does not guarantee meaningful connection, it provides the structural conditions that support the kinds of relational, embodied, and playful interactions participants described.
The study underscores that youth are not necessarily seeking less time on social media, but rather more fulfilling and emotionally present interactions within these digital spaces. Their visions point towards a future where social media feels more like a dynamic, shared social world like those found in 3D gaming environments. While our findings are rooted in the experiences of youth already familiar with 3D gaming, they offer a hopeful design direction for spaces that support meaningful social connections that youth desire.

### v3

@@ 4.5.2. Making Social Media About (Meaningful) Connection, Not (Passive) Consumption
Beyond content navigation, participants emphasized shifting the platform’s core purpose from endless consumption toward meaningful social connection. They expressed frustration with current platforms promoting mindless, passive consumption and hoped SMH would foster meaningful, intentional connections through active social engagement. They described frustration with how easily time slips away on platforms like Instagram and TikTok, leaving them unfulfilled. P14 captured this sentiment, describing spending time on Instagram Reels while “literally just laying in [their] bed” as “wasting [their] time.”
MSM’s infinite scrolling design contributed to this feeling. P11 explained “there’s always more content on social media,” fueling fear of missing out and urges to keep scrolling. P14 critiqued TikTok for exposing users to “clickbaity” low-quality content that “doesn’t stimulate [users] at all” and “messes with [their] attention span.”
In contrast, participants described active, social engagement as far more rewarding. P16 shared that “anytime not scrolling” but “talking to people” feels worthwhile, “even if it’s just light conversation.” P14 noted meaningful use comes when “actually talking to someone,” especially staying connected with people “not in close proximity.” They reflected that while users might spend more time on SMH, it would feel worthwhile because “at the end of the day, you’re connecting with a person that you can’t otherwise connect with… and that makes it more valuable in and of itself” compared to watching “50-second TikTok videos of somebody making slime for like an hour.”
Participants contrasted passive consumption with active participation in interactive environments. P10 explained that while doomscrolling feels unproductive, activities like Minecraft offer “progress” that can be “saved and come back to.” P06 noted they “don’t feel bad… most of the time [they] use Roblox,” because it involves “playing games with friends” and “customizing spaces.”
Reflecting on SMH’s potential, P09 acknowledged “time spent might not be mitigated,” but emphasized “the way you’re spending it” would make the difference. They explained SMH would “feel a lot more interactive instead of just taking in content, and more and more and more content,” suggesting meaningful engagement, rather than passive consumption, ultimately mattered.
@@ 4.6. Playfulness and Whimsy
Participants described SMH as a platform feeling youthful, playful, and deeply immersive, offering entertainment plus opportunities for creative self-expression, social bonding, and emotional well-being. By embracing whimsy and delight, SMH would make social media feel more personal, less pressured, and more meaningful.
@@ 4.6.1. Embracing Creative Joy and Playful Self-Expression
Participants imagined SMH transforming social media into a space for creative joy and playful self-expression, offering freedom to experiment with identity, explore personal interests, and connect through creativity. P11 imagined a “dream social media world” where users could fully redesign their profiles, “maybe their avatar in the middle, maybe their posts on either side, maybe a bio below”, allowing them to “express their creativity and, like, yourself” in new ways not possible on conventional platforms.
For some, personal customization felt like meaningful identity exploration. P01 described the deeper satisfaction in “just that process of creating it, and finding out who you are through that process.” P21 echoed this sentiment, describing SMH as feeling “more like a game” because of the “creative freedom” it provides, making the experience feel “more fictional than real” in a way that encourages experimentation and self-expression. P22 added that SMH offered something they hadn’t found anywhere else, explaining, “I can really be creative here… I can make something that I haven’t made anywhere else, you know. That’s where the fun aspect comes in, really, for me.”
This creativity extends beyond individuals to the wider environment and community. Participants described the excitement of exploring what others had created, likening it to visiting realms in Minecraft where you “get to see how like everyone puts in their own effort into it… Whoa! They built that, that’s so cool” (P06). SMH felt alive and engaging whether users were socializing or spending time alone: “it’s not just like, Oh, the group chats are only fun when there’s other people here. No, it’s fun when we’re by ourselves, too, we get to do our own things” (P06). The interactive environment itself, where “you’re not just sitting there chatting, you’re also walking around the space… you’re interacting with whatever else is in that area” (P06), offered more to do than conventional platforms. Whether through building, exploring, or engaging with shared content, SMH transformed social interaction into something playful, creative, and deeply engaging.
@@ 4.6.2. Designing Engaging and Playful Privacy Mechanisms
Participants imagined privacy not as tedious and confusing but as part of playful experience, envisioning interactive, game-like mechanisms turning boundary-setting into something fun, discoverable, and socially meaningful. P11 envisioned controlling space access through gestures like “secret knock” or hidden objects revealing “secret passages” for trusted friends. Others suggested knowledge-based gates like “secret handshakes” (P21) shared among friends. Drawing from the Marauder’s Map, P02 imagined privacy controls activated by “spells”, unlocking spaces when correctly summoned.
P05 compared these features to “little Easter eggs,” remarking “there’s that little kid part of everyone” that makes such features enjoyable, contrasting them with mainstream social media’s “wall of settings” that often feel confusing.
@@ 4.6.3. Augmenting Everyday Interactions with Fantastical Interfaces and Environments
Participants imagined SMH using magical and fantastical technologies to transform everyday digital interactions, making them feel more immersive, accessible, and emotionally uplifting, even for those with limited real-world opportunities. P01 questioned the need for phones at all, envisioning wands projecting “holographic screens, like Jarvis in Iron Man,” or activating interfaces through spells. Participants also imagined alternative magical devices: VR glasses, a “makeup mirror” portal, or a “watch” that immerses users into SMH.
@@ 4.7. Expanding Horizons
Given the speculative FI co-design workshop, SMH embodies youth aspirations and values for social media platform interactions. While acknowledging potential challenges like over-immersion or distraction, they viewed these downsides as worthwhile trade-offs for its benefits. Participants described SMH as an inspiring alternative to existing social media, offering self-expression opportunities and deeply personal, intentional connection. They saw SMH not just as a better platform, but as an inspiring redefinition of social media’s purpose and potential.
@@ 4.7.1. Breaking Free from Platform Monotony with Novel Interactions
Participants saw SMH as offering better user experiences, departing from existing platforms’ repetitive structures, and inspiring hope for social media’s future. P06 appreciated how SMH combines desirable aspects: “You get friends who you can talk to, keep up with, and stay connected to. And you get things you can customize. You kind of get the whole package there.”
Participants appreciated SMH’s break from existing platform monotony. P11 pointed out that most social media platforms share a common “standard” with repetitive features like “feed, messages, settings,… profile.” In contrast, SMH introduced new types of interactions that deviated from familiar patterns. P02 described SMH as “a combination of Discord and BeReal” that brings together casual sharing and topic-based discussions. P17 envisioned SMH as “more interactive, kind of like if Discord and Instagram were combined,” enjoying Discord’s interactions with Instagram’s profiles and posting ability. P22 highlighted limited customization on other platforms, explaining “I just have a profile picture.” On SMH, they could “really be creative” and “make something that [they] haven’t made anywhere else,” making the experience feel “novel,” like a “breath of fresh air.”
For some, SMH’s novelty lay in prioritizing enjoyment over seriousness. P14 described initially defaulting to the idea of a more “mature” 2D version, but later embracing the playful, immersive 3D version as more personally meaningful and fun. For example, P14 initially preferred a 2D version but shifted to 3D, reasoning:
@@
P22 described SMH as “the coolest,” “childlike,” “less generic,” and “a lot more fun” compared to other platforms. P06 observed that “the physical space” on SMH felt “a lot more fantastical… whimsical, fun” compared to “traditional ones.” The platform also evokes nostalgia and happiness. P15 remarked “Having this app would actually make them feel happy… with nostalgia,” anticipating SMH would be “a super big hit” with teenagers due to its playful, engaging nature.
@@ 4.7.2. Inspiring Hope and Expanding the Scope of Social Media Interactions
Through the FI process, participants realized they weren’t limited by existing platform structures. Instead, they felt empowered to imagine social technologies centering joy, creativity, and authentic human connection–qualities they had once believed beyond digital spaces’ reach. P10 pointed out that overcoming the design fixation around social media is crucial:
@@ 5. Discussion
Our analysis shows that youth imagine future social media not as a rejection of current platforms but as a reconfiguration of their strengths. They want the visibility and reach of social networking and broadcasting platforms while also craving the intimacy and authenticity found in spaces like games, voice chat, or small-group apps. Figure 2 illustrates this blended vision: an aspirational middle ground where meaningful connection is not an afterthought but the central purpose. Building on this insight, the following subsections trace how participants moved from critiquing existing platforms toward articulating design goals, culminating in our proposal of “spatial integrity” as a framework for supporting deeper forms of online connection.
The relevance of these findings has become increasingly apparent with recent technological developments. With the introduction of technologies like Genie 3 and similar platforms, the fictional world medium of remote social connection that participants envisioned is barely fictional. The 3D, open-world, and fully customizable platforms for self-expression and organic social connection that emerged from our workshops now represent concrete possibilities for the near future, much nearer than we anticipated when we began this study. This technological trajectory underscores the timeliness of our findings and their potential to guide the development of relationship-supporting technologies that leverage these emerging capabilities (DeepMind, 2025).
5.1. From Reactionary Designs to Aspirational Design Goals
However, taking social media away from young people or framing it as inherently harmful oversimplifies a much more complex reality. Despite public concern about social media’s role in youth mental health, the causal evidence remains inconclusive, and youth themselves report a mix of positive and negative experiences (Vogels, 2023). At the same time, research shows that fear-based policies and alarmist narratives can do harm of their own, amplifying anxiety, mistrust, and feelings of helplessness among youth, rather than supporting constructive engagement (Kim et al., 2024a; Lee and Hancock, 2024). This misalignment suggests that the solution lies not in rejecting social media altogether, but in asking what these platforms could become if they were designed to better reflect the kinds of social connections that youth actually seek. Rather than accepting the limitations of existing platforms as inevitable, there is an opportunity to rethink their underlying priorities and architectures.
However, in our earlier work with this youth population, we noticed several recurring challenges that shaped how participants thought about social media. First, there was a strong design fixation on social media design around existing apps, both in terms of interface, such as small 2D mobile screens, endless scrolling, and uniform templates, and in terms of user experience, where visibility depends on maintaining large networks of friends and followers, often at the expense of authenticity and comfort. Second, participants showed a cognitive fixation when imagining the future of social media, struggling to see it as anything beyond its currently perceived role as a convenient but limited tool for remote social connection. Third, these fixations were compounded by generally negative experiences with social media, which led many in earlier interviews to propose reactionary fixes, like stronger privacy, the removal of infinite scrolls, or eliminating general media content such as Reels. Such cognitive fixation not only overlooks how social features can be designed differently—as demonstrated by their own designs—but also reinforces negative perceptions of social media that prior literature suggests can impact psychological well-being and quality of life (Lee and Hancock, 2024; Kim et al., 2024a).
To help youth envision beyond these constraints, we leveraged the FI method, immersing participants in a magical world setting. The procedure was intended to serve three purposes: breaking participants free from these biases, learned helplessness, and pessimism by encouraging them to think beyond the “app” or “platform” on mobile devices; equipping them with prompts that licensed thinking beyond perceived technical and practical constraints; and creating conditions where they could explore proactive and aspirational visions of social interaction rather than simply reacting to existing frustrations. Our chosen method prioritized uncovering underlying desires and needs that young people associate with meaningful remote social interaction over producing concrete feature lists.
In line with the goals of co-design processes (Sanders and Stappers, 2008), our interest lay less in the specifics of functionality and more in uncovering what kinds of interactions and environments participants valued most. Particularly noteworthy was the fact that the workshops helped participants expand their understanding of what social platforms could be, empowering them to feel more hopeful and articulate desires they might have previously dismissed as impossible within conventional social media paradigms. This demonstrates how the design method chosen in our study offered a generative path toward hope in domains where narratives of doom are prevalent.
Our analysis contributes insights about the types of remote social interactions that youth find most meaningful and fulfilling, examining how these desired interactions align with or diverge from current platform affordances, as we detail in Section 5.2. We also identify core aspects of design that can undermine relational quality from which we develop the “spatial integrity” framework introduced in Section 5.3. Taken together, we hope this work equips designers with proactive design goals for envisioning new futures of social technologies and provides an initial framework for evaluating design integrity in relationship-supporting aspects.
@@ 5.4. Limitation and Future Work
Our study has several key limitations to consider. First, all participants had prior experience with 3D gaming platforms, so the findings do not inherently transfer to youth who are not familiar with such environments. Second, although we were cognizant of how fixed perceptions around “social media” might constrain ideation, our continued use of this terminology throughout the interviews may have inadvertently reinforced existing mental models rather than encouraging completely novel conceptualizations. Third, while the Harry Potter narrative framework successfully transported participants into a magical design space, our specific scenario about a Muggle-born student introducing social technology sometimes led participants to focus on wizard-Muggle divisions rather than the intended focus on connection across distance. This suggests opportunities to refine the narrative approach.
@@ 6. Conclusion
This study explored how youth with 3D gaming experience reimagine social media when freed from current platform constraints. Through Hogwarts’ fictional world, participants envisioned social experiences that felt spatial, relational, and emotionally meaningful, contrasting the passive, feed-based consumption defining mainstream social media.
Their visions called for platforms that feel more like 3D games than Instagram, emphasizing intentional navigation, ambient social presence, customizable spaces, and playful privacy mechanisms that adapt to real relationships’ complexity while incorporating playfulness and whimsy.
Building on these insights, we introduced spatial integrity, foundational design principles enabling such visions. While spatial integrity alone cannot guarantee meaningful connection, it provides structural conditions supporting the relational, embodied, and playful interactions participants described.
The study underscores that youth seek not less time on social media, but more fulfilling and emotionally present interactions within digital spaces. Their visions point toward a future where social media feels like the dynamic, shared social worlds found in 3D gaming environments. While our findings stem from youth familiar with 3D gaming, they offer a hopeful design direction for spaces supporting the meaningful social connections youth desire.